# The Low-Degree Hardness of Finding Large Independent Sets in Sparse Random Hypergraphs

### v2

@@ 4.3 Forbidden Structures
$|S_{i}|\,\leq\,s\,\gamma_{i}\,r\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}.$
Then for $p=d/n^{r-1}$ there is no independent set $S^{\prime}$ in $\mathcal{H}(r,n,p)$ satisfying
$|S^{\prime}\cap S_{i}|\,\geq\,\varepsilon\,\gamma_{i}\,r\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r,$
@@
Without loss of generality, assume $i^{\star}=r$. Let $\Phi$ be as defined in the proof of Proposition 4.4. We will compute the expected number of such sets $S^{\prime}$ and show it is exponentially small (the proof then follows by Markov’s inequality). Let $N$ denote the number of possible $S^{\prime}$. We have
$\mathbb{E}[N]$ $\leq\left(\prod_{i=1}^{r}\binom{|S_{i}|}{\varepsilon\gamma_{i}r\Phi}\right)\left(1-\frac{d}{n^{r-1}}\right)^{\varepsilon^{r}\,r^{r}\,\Phi^{r}\prod_{i=1}^{r}\gamma_{i}}$
$\leq\left(\prod_{i=1}^{r}\binom{s\,\gamma_{i}\,r\Phi}{\varepsilon\gamma_{i}r\Phi}\right)\exp\left(-\frac{d}{n^{r-1}}\,\varepsilon^{r}\,r^{r}\,\Phi^{r}\prod_{i=1}^{r}\gamma_{i}\right)$
@@
$=\exp\left(-\Phi\left(\frac{\varepsilon^{r}\,\gamma_{r}r\,\log d}{r-1}-\varepsilon\,r\,\log\left(\frac{es}{\varepsilon}\right)\right)\right)$
@@ 4.4 Proof of the Intractability Result
For any $r\geq 2$, $\varepsilon>0$, and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)$, $\sum_{i=1}^{r}\gamma_{i}=1$, and $i^{\star}=\arg\max_{i}\gamma_{i}$, there exist $K,\,d_{0}>0$ such that for any $d\geq d_{0}$ there exists $n_{0},\eta,C_{1},C_{2}>0$ such that for any $n\geq n_{0}$, $\xi\geq 1$, $1\leq D\leq\dfrac{C_{1}n}{\xi\log n}$, and $\delta\leq\exp\left(-C_{2}\xi\,D\log n\right)$, if
$k_{i}\,\geq\,(1+\varepsilon)\,\gamma_{i}\,r\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r,$
@@
$\Phi=\left(\frac{\gamma_{r}\,\log d}{d\,r^{r-1}\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\qquad\tilde{\varepsilon}=\varepsilon\min_{i\in[r]}\gamma_{i}.$
@@
Suppose for contradiction there exists a $(D,K-1,c)$-stable algorithm $f$. Sample the interpolation path $A^{(0)},\ldots,A^{(T)}$ as in Definition 2.2 and let $I_{t}=V_{f}^{\eta}(A^{(t)})$ be the resulting independent sets for $0\leq t\leq T$. We construct a sequence of sets $S_{1},\ldots,S_{K}$ as follows: let $S_{1}=I_{0}$, and for $k\geq 2$, let $S_{k}$ be the first $I_{t}$ such that $|I_{t}\setminus\bigcup_{j<k}S_{j}|\geq\tilde{\varepsilon}\,n\,\Phi/4$ (if no such $t$ exists, we say the process fails). We define the following events:
@@
The following claim shows that if the events (EB1) and (EB2) occur, then so does (EB3).
@@
Conditioned on such an outcome, we will show that the procedure to define $S_{k}$ succeeds with high probability. Suppose for some $0\leq T^{\prime}\leq T-m$, we have sampled $A^{(0)},\ldots,A^{(T^{\prime})}$ and successfully selected $S_{1}=I_{0},\ldots,S_{K^{\prime}}=I_{t_{K^{\prime}}}$ for some $1\leq K^{\prime}<K$. Let $S=\cup_{k\leq K^{\prime}}S_{k}$ and note that $A^{(T^{\prime}+m)}$ is independent of $A^{(T^{\prime})}$. Applying Lemma 4.6 with
@@
In particular, $S_{K^{\prime}+1}=I_{t}$ for some $T^{\prime}<t\leq T^{\prime}+m$ and thus the process succeeds by timestep $T=(K-1)m$ with high probability. We may conclude by a union bound over $t$ that (EB1) occurs with probability at least $1-(T+1)\delta-\exp\left(-\Omega(n)\right)$.
Since $f$ is assumed to be $(D,K-1,c)$-stable, (EB2) occurs with probability at least $p^{4(K-1)D/c}$, where $p=d/n^{r-1}$. By Proposition 4.4, we may conclude that at least 1 of the events (EB1)(EB3) occurs with probability at most
@@
It is now enough to show that the above is $<1$, which is implied by the following for some constant $C>0$ sufficiently large:
@@ Acknowledgements
We thank Will Perkins and Cheng Mao for helpful comments.
[…]

### v3

@@ 4.3 Forbidden Structures
$|S_{i}|\,\leq\,s\,\gamma_{i}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}.$
Then for $p=d/n^{r-1}$ there is no $\bm{\gamma}$-balanced independent set $S^{\prime}$ in $\mathcal{H}(r,n,p)$ satisfying
$|S^{\prime}\cap S_{i}|\,\geq\,\varepsilon\,\gamma_{i}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r,$
@@
Without loss of generality, assume $i^{\star}=r$. Let $\Phi$ be as defined in the proof of Proposition 4.4. The probability that such a set $S^{\prime}$ exists is at most
$\left(\prod_{i=1}^{r}\binom{|S_{i}|}{\varepsilon\gamma_{i}r\Phi}\right)\left(1-\frac{d}{n^{r-1}}\right)^{\varepsilon^{r}\,r^{r}\,\Phi^{r}\prod_{i=1}^{r}\gamma_{i}}$ $\leq\left(\prod_{i=1}^{r}\binom{s\,\gamma_{i}\,r\Phi}{\varepsilon\gamma_{i}r\Phi}\right)\exp\left(-\frac{d}{n^{r-1}}\,\varepsilon^{r}\,r^{r}\,\Phi^{r}\prod_{i=1}^{r}\gamma_{i}\right)$
@@
$=\exp\left(-r\Phi\left(\frac{\varepsilon^{r}\,\gamma_{r}\,\log d}{r-1}-\varepsilon\,\log\left(\frac{es}{\varepsilon}\right)\right)\right)$
@@ 4.4 Proof of the Intractability Result
For any $r\geq 2$, $\varepsilon>0$, and $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ such that $\gamma_{i}\in(0,1)\cap\mathbb{Q}$, $\sum_{i=1}^{r}\gamma_{i}=1$, and $i^{\star}=\arg\max_{i}\gamma_{i}$, there exist $K,\,d_{0}>0$ such that for any $d\geq d_{0}$ there exists $n_{0},\eta,C_{1},C_{2}>0$ such that for any $n\geq n_{0}$, $\xi\geq 1$, $1\leq D\leq\dfrac{C_{1}n}{\xi\log n}$, and $\delta\leq\exp\left(-C_{2}\xi\,D\log n\right)$, if
$k_{i}\,\geq\,(1+\varepsilon)\,\gamma_{i}\,n\left(\frac{\gamma_{i^{\star}}\log d}{d\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\quad\text{for each }1\leq i\leq r,$
@@
$\Phi=\frac{1}{r}\left(\frac{\gamma_{r}\,\log d}{d\,(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)},\qquad\tilde{\varepsilon}=\varepsilon\min_{i\in[r]}\gamma_{i}.$
@@
Suppose for contradiction there exists a (D,K1,c)(D,K-1,c)-stable algorithm $f$. Sample the interpolation path $A^{(0)},\ldots,A^{(T)}$ as in Definition 2.3 and let $I_{t}=V_{f}^{\eta}(A^{(t)})$ be the resulting independent sets for $0\leq t\leq T$. We construct a sequence of sets $S_{1},\ldots,S_{K}$ as follows: let $S_{1}=I_{0}$, and for $k\geq 2$, let $S_{k}$ be the first $I_{t}$ such that $|I_{t}\setminus\bigcup_{j<k}S_{j}|\geq\tilde{\varepsilon}\,n\,\Phi/4$ (if no such $t$ exists, we say the process fails). We define the following events:
@@
The following claim shows that if the events (EB1) and (EB2) occur, then (EB3) does not.
@@
On the event that this occurs, we will show that the procedure to define $S_{k}$ succeeds with high probability. Suppose for some $0\leq T^{\prime}\leq T-m$, we have sampled $A^{(0)},\ldots,A^{(T^{\prime})}$ and successfully selected $S_{1}=I_{0},\ldots,S_{K^{\prime}}=I_{t_{K^{\prime}}}$ for some $1\leq K^{\prime}<K$. Let $S=\cup_{k\leq K^{\prime}}S_{k}$ and note that $A^{(T^{\prime}+m)}$ is independent of $A^{(T^{\prime})}$. Applying Lemma 4.6 with
@@
In particular, $S_{K^{\prime}+1}=I_{t}$ for some $T^{\prime}<t\leq T^{\prime}+m$ and thus the process succeeds by timestep T=(K1)mT=(K-1)m with high probability. We may conclude by a union bound over $t$ that (EB1) occurs with probability at least $1-(T+1)\delta-\exp\left(-\Omega(n)\right)$.
Since $f$ is assumed to be (D,K1,c)(D,K-1,c)-stable, (EB2) occurs with probability at least $p^{4(K-1)D/c}$, where $p=d/n^{r-1}$. By Proposition 4.4, we may conclude that at least 1 of the events (EB1)(EB3) occurs with probability at most
@@
It is now enough to show that the above is <1<1, which is implied by the following for some constant $C>0$ sufficiently large:
@@ Acknowledgments
We thank Will Perkins and Cheng Mao for helpful comments. We also thank the anonymous referee for carefully reading this manuscript and providing helpful suggestions.
Appendix A Sparse Random Hypergraphs are Locally Hypertree-Like
In this section, we give the proof of our claim that $\mathcal{H}_{r}(n,p)$ is locally tree-like for $p=d/\binom{n-1}{r-1}$ (this fact was used in the proof of Lemma 2.1). Our proof is nearly identical to that of [Bor16, Theorem 3.12] for graphs. There are two key lemmas in their proof that differ in the hypergraph setting (albeit slightly). We include the proofs of these lemmas in this section.
We begin by showing that $\mathcal{H}_{r}(n,p)$ “looks like” $\mathbb{T}_{d}^{GW}$ locally. The essence of “looks like” in the context of local weak convergence is to show that the $s$-neighborhood of a random vertex in $\mathcal{H}_{r}(n,p)$ has a distribution that converges to the $s$-level subtree of $\mathbb{T}_{d}^{GW}$ as $n\to\infty$. More formally, we prove the following proposition:
Proposition A.1.
Fix any integers $d,r,s>0$ and $\varepsilon>0$. Then for sufficiently large $n$ and $p=d/{n-1\choose r-1}$, we have
$\mathbb{P}_{H\sim\mathcal{H}_{r}(n,p)}\left[\left|\{u\in V(H)\,:\,B_{H}(u,s)\ncong\mathbb{T}_{d}^{GW}\}\right|\geq\varepsilon n\right]\leq\varepsilon.$
Consider a hypergraph $H=(V,E)$ and let $v\in V$ be smallest vertex under some arbitrary order. The proof follows the breadth-first search (BFS) of the connected component $H(v)\subseteq H$ containing $v$.
This exploration is iterative. At integer step $t$, a vertex may belong to the active set $A_{t}$, to the unexplored set $U_{t}$, or to the connected set $C_{t}=V\setminus(A_{t}\cup U_{t})$. We begin with $A_{0}=\{v\}$, $C_{0}=\emptyset$, and $U_{0}=V\setminus\{v\}$. For integer $t\geq 0$, if $A_{t}\neq\emptyset$, we define $v_{t+1}$ as the vertex in $A_{t}$ with the minimum order. Let $E_{t+1}=\{e\in E\,:\,v_{t+1}\in e,\,e\setminus\{v_{t+1}\}\subseteq U_{t}\}$ be the edges incident to $v_{t+1}$ that have not yet been explored in the BFS, and let $I_{t}=\{v\in U_{t}\,:\,v\in e\text{ for some }e\in E_{t+1}\}$ be the neighbors of $v_{t+1}$ in $U_{t}$. We set
$\begin{cases}A_{t+1}=(A_{t}\cup I_{t+1})\setminus\{v_{t+1}\}\\ U_{t+1}=U_{t}\setminus I_{t+1}\\ C_{t+1}=C_{t}\cup\{v_{t+1}\}\end{cases}$
The process stops when $A_{t}=\emptyset$. It follows by construction that $|H(v)|=\inf\{t\geq 1:A_{t}=\emptyset\}$.
For ease of notation, we set $X_{t+1}=|E_{t+1}|$ and $\tau=|H(v)|$ so that for $t<\tau$, we have
$|A_{t}|\leq 1+\sum_{s=1}^{t}((r-1)X_{s}-1),\qquad|U_{t}|=n-t-|A_{t}|,\qquad|C_{t}|=t.$
Now, we consider the BFS with starting vertex $v$ and the graph $H\sim\mathcal{H}_{r}(n,p)$. We define the filtration $\mathcal{F}_{t}=\sigma((A_{0},U_{0},C_{0}),\dots,(A_{t},U_{t},C_{t}))$. The hitting time $\tau\coloneqq\inf\{t\geq 1:A_{t}=\emptyset\}$ is a stopping time for this filtration. Notice that for any integer $t\geq 0$, given $\mathcal{F}_{t}$, if $\{t<\tau\}\in\mathcal{F}_{t}$ then $X_{t+1}\sim{\rm Bin}\left({|U_{t}|\choose r-1},d/{n-1\choose r-1}\right)$.
With this set-up in hand, we first show the convergence of the exploration.
Lemma A.2.
On an enlarged probability space, there exists a sequence $(X_{t}^{\prime})_{t\geq 1}$ of i.i.d. ${\rm Poiss}(d)$ variables such that
$\mathbb{P}\left[\left(X_{1},\dots,X_{t\wedge\tau}\right)\neq\left(X_{1}^{\prime},\dots,X_{t\wedge\tau}^{\prime}\right)\right]\leq\frac{rd(1+rd)t^{2}+td}{n-1}.$
Proof.
The stopping property implies that $\{t<\tau\}$ is $\mathcal{F}_{t}$-measurable. Note that if $\{t<\tau\}$ holds, then
$\mathbb{E}(X_{t+1}|\mathcal{F}_{t})=\frac{d{|U_{t}|\choose r-1}}{{n-1\choose r-1}}\leq d.$
In particular, we have
$\mathbb{E}\left[(|A_{t}|-1)\mathbbm{1}_{t<\tau}\right]\leq\sum_{s=0}^{t-1}\mathbb{E}[((r-1)X_{s+1}-1)\mathbbm{1}_{s<\tau}]\leq rdt.$ (13)
Now, on an enlarged probability space, let $\xi_{t+1}$ be, given $\mathcal{F}_{t}$, a binomial variable ${\rm Bin}\left({n-1\choose r-1}-{|U_{t}|\choose r-1},d/{n-1\choose r-1}\right)$ independent of $X_{t+1}$. Then $Y_{t+1}=X_{t+1}+\xi_{t+1}$ is a binomial variable ${\rm Bin}\left({n-1\choose r-1},d/{n-1\choose r-1}\right)$ and $(Y_{t})_{t\geq 1}$ is an i.i.d. sequence. By a union bound, we have
$\mathbb{P}\left[\left(X_{1},\dots,X_{t\wedge\tau}\right)\neq\left(Y_{1},\dots,Y_{t\wedge\tau}\right)\right]\leq\sum_{s=1}^{t}\mathbb{E}\left[\mathbbm{1}_{s<\tau}\mathbb{P}\left[X_{s}\neq Y_{s}|\mathcal{F}_{s}\right]\right]\leq\sum_{s=1}^{t}\mathbb{E}\left[\mathbbm{1}_{s<\tau}\mathbb{P}\left[\xi_{s}\neq 0|\mathcal{F}_{s}\right]\right].$
It is straightforward to compute the probability that $\xi_{s}\neq 0$:
$\mathbb{P}\left[\xi_{s}\neq 0|\mathcal{F}_{s}\right]$ $=1-\left(1-\frac{d}{{n-1\choose r-1}}\right)^{{n-1\choose r-1}-{|U_{s}|\choose r-1}}$
$\leq\frac{d}{{n-1\choose r-1}}\left({n-1\choose r-1}-{|U_{s}|\choose r-1}\right)$
$=\frac{d}{{n-1\choose r-1}}\left({n-1\choose r-1}-{n-t-|A_{s}|\choose r-1}\right)~~~~~~~~~~(\text{since }|U_{s}|=n-s-|A_{s}|)$
$\leq\frac{d}{{n-1\choose r-1}}\cdot(s+|A_{s}|-1){n-2\choose r-2}~~~~~~~~~~\left(\text{using }{n\choose r}-{n-k\choose r}\leq k{n-1\choose r-1}\right)$
$\leq\frac{rd(t+|A_{t}|-1)}{n-1},$
where we use the fact that $s\leq t$ and $A_{s}\subseteq A_{t}$. By (13), we have
$\mathbb{P}\left[\left(X_{1},\dots,X_{t\wedge\tau}\right)\neq\left(Y_{1},\dots,Y_{t\wedge\tau}\right)\right]\leq\sum_{s=1}^{t}\frac{rd(t+rdt)}{n-1}\leq\frac{rd(1+rd)t^{2}}{n-1}.$ (14)
We use the following simple bound on the total variation distance between binomial and Poisson random variables:
$d_{TV}\left({\rm Bin}\left({n-1\choose r-1},d/{n-1\choose r-1}\right),{\rm Poiss}(d)\right)\leq\frac{d}{{n-1\choose r-1}}$,
which implies
$d_{TV}(\left(Y_{1},\dots,Y_{t}\right),{\rm Poiss}(d)^{\otimes t})\leq\frac{td}{{n-1\choose r-1}}\leq\frac{td}{n-1},$ (15)
for $r\geq 2$. Combining (14) and (15), the claim follows by the maximal coupling inequality. ∎
Lemma A.3.
For integer $t\geq 0$, let $J_{t}=\{(u_{1},\cdots,u_{r-1})\subset A_{t}:(v_{t+1},u_{1},\cdots,u_{r-1})\in E\}$. We have that
$\mathbb{P}[\exists 1\leq s\leq t\wedge\tau:|J_{s}|\neq 0]\leq\frac{rd^{2}t^{2}}{n-1}$
Proof.
Given $\mathcal{F}_{t}$, if $t<\tau$, $|J_{t}|$ is a binomial random variable ${\rm Bin}\left({|A_{t}|-1\choose r-1},d/{n-1\choose r-1}\right)$. The union bound yields
$\mathbb{P}\left[\exists 1\leq s\leq t\wedge\tau:|J_{s}|\neq 0\right]\leq\sum_{s=1}^{t}\mathbb{E}\left[\mathbbm{1}_{s<\tau}\mathbb{P}[|J_{s}|\neq 0|\mathcal{F}_{s}]\right].$
Since
$\mathbb{P}[|J_{s}|\neq 0|\mathcal{F}_{s}]=1-\left(1-\frac{d}{{n-1\choose r-1}}\right)^{{|A_{s}|-1\choose r-1}}\leq\frac{d(|A_{s}|-1)}{n-1},$
we have
$\mathbb{P}\left[\exists 1\leq s\leq t\wedge\tau:|J_{s}|\neq 0\right]\leq\frac{d}{n-1}\sum_{s=1}^{t}\mathbb{E}\left[(|A_{s}|-1)\mathbbm{1}_{s<\tau}\right].$
Using the bound (13) again gives
$\mathbb{P}\left[\exists 1\leq s\leq t\wedge\tau:|J_{s}|\neq 0\right]\leq\frac{d}{n-1}\cdot rdt^{2}=\frac{rd^{2}t^{2}}{n-1},$
as desired. ∎
With these lemmas in hand, the proof of Proposition A.1 is identical to the graph case, mutatis mutandis. One can, for example, follow the strategy of Theorem 3.12 in the lecture notes of Bordenave [Bor16] by replacing Lemmas 3.13 and 3.14 with Lemmas A.2 and A.3, respectively.